# Energy-Efficient Balanced Flow Control Achieved through Optimization of Synthetic Jet Placement Based on Deep Reinforcement Learning

### v1

@@ 3.2.2 Scenario B: Energy Consumption with Simultaneous Control (Two Pairs)
Figure 12: Comparison of action values over non-dimensional time for different synthetic jets configurations around the cylinder. (a) Two pairs of jets: $J_{3}^{1},J_{3}^{2}$ and $J_{4}^{1},J_{4}^{2}$. (b) Two pairs of jets: $J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$.
We execute flow control with a strategy involving the mass flow rates of two pairs of synthetic jets, referred to as Action in figure 12. The action values, representing the external energy consumed, are plotted over non-dimensional time for various jet configurations around the cylinder. Figure 12(a) depicts the control strategy utilizing two pairs of synthetic jets ($J_{3}^{1},J_{3}^{2}$ and $J_{4}^{1},J_{4}^{2}$). The initial peak in action values indicates the maximum energy required at the onset of the flow control process. As time progresses, the energy consumption stabilizes, suggesting that the control strategy becomes more consistent. We observe that the jet pair $J_{4}^{1},J_{4}^{2}$ consumes slightly less energy in the long run compared to $J_{3}^{1},J_{3}^{2}$, indicating that positioning synthetic jets at $J_{4}$ (75°) may offer an energy-saving advantage over positioning at $J_{3}$ (90°).
The energy consumption for the control strategy utilizing two pairs of synthetic jets ($J_{3}^{1},J_{3}^{2}$ and $J_{2}^{1},J_{2}^{2}$) is shown in figure 12(b). The initial action peak is again prominent, reflecting the energy required at the start of the flow control process. Over time, both configurations exhibit reduced energy consumption, with action values stabilizing. However, jet pair $J_{3}^{1},J_{3}^{2}$ demonstrates a slightly higher energy demand over time compared to $J_{2}^{1},J_{2}^{2}$, suggesting that $J_{2}^{1},J_{2}^{2}$ may be more energy-efficient in maintaining effective flow control. This analysis provides a deeper understanding of the energy efficiency of different jet configurations in controlling flow around the cylinder. The initial peaks highlight the energy demands at the start, while the subsequent stabilization reflects the sustained maintenance of flow control. The differences in energy consumption among configurations can guide the selection of jet positions to optimize flow control efficiency, with the $J_{4}$ position (75°) and $J_{2}$ position (105°) showing potential as a more energy-efficient choice.
@@ 3.2.3 Scenario B: Energy Consumption with Simultaneous Control (Three Pairs)
A detailed analysis of energy consumption during flow control using three pairs of synthetic jets is presented in figure 13. These jets ($J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$) are symmetrically arranged around the cylinder at angular positions of 75° ($J_{4}$), 90° ($J_{3}$), and 105° ($J_{2}$) relative to the forward stagnation point (0°). The action values shown in figure 13 correspond to the external energy expenditure associated with jet operation over non-dimensional time. In this context, $E_{J_{3}}$ denotes the external energy required by the synthetic jets pair $J_{3}$, comprising $J^{1}_{3}$ and $J^{2}_{3}$, to achieve effective flow control. Interestingly, when compared to figure 12, where only two pairs of synthetic jets are employed, figure 13 demonstrates that the activation of three pairs of synthetic jets leads to a reduction in the peak external energy demand for each individual jet. This suggests that the simultaneous use of additional jets can distribute the energy load more effectively, thereby enhancing the overall efficiency of the flow control strategy.
@@
Figure 13: Energy consumption analysis for different jet pairs positioned around a cylinder. The left diagram illustrates the angular positions of the jet pairs $J_{2}^{1},J_{3}^{1},J_{4}^{1}$ on the upper half and $J_{2}^{2},J_{3}^{2},J_{4}^{2}$ on the lower half, with angles ranging from 75° to 105°. The right graph displays the corresponding action values over non-dimensional time for each jet configuration.
At the onset, the action values across all three configurations, as optimized via DRL, exhibit pronounced peaks. Following the initial transient phase, the energy consumption stabilizes across all configurations, indicating a uniform energy requirement as the flow control reaches a steady state. The jet pair positioned at 90° ($J_{3}$) demonstrates the highest peak, suggesting that the energy consumption at the 90° jet position ($E_{J_{3}}$) is significantly higher than that of the other two configurations ($E_{J_{4}}\approx E_{J_{2}}$). This observation implies that the 90° jet position demands greater energy to achieve comparable control efficacy, thereby indicating a lower energy efficiency compared to the 75° and 105° positions. The variations in energy consumption associated with different jet positions for flow control around a cylinder, emphasizing the critical importance of selecting an optimal jet position to minimize energy consumption while maintaining effective flow regulation, are highlighted in figure 13. The results suggest that the jet positions at 75° ($J_{4}$) and 105° ($J_{2}$) exhibit similar and reduced energy demands, indicating their superiority in sustaining flow control. Conversely, although the 90° position is frequently utilized in flow control applications, it may not represent the most energy-efficient choice. These findings are vital for refining flow control strategies, particularly in scenarios where energy efficiency is of paramount importance.
@@ 4.1 A pair of synthetic jets controllers
4.1.1 Scenario A: Control Performance of Jet Pair
Figure 14: Performance evaluation of different jet configurations for flow control around a square cylinder. (a) Reward curves over training episodes for various jet positions. (b) Drag coefficient $C_{D}$ over non-dimensional time, comparing the baseline scenario with different jets configurations. (c) Lift coefficient $C_{L}$ over non-dimensional time, comparing the baseline scenario with different jets configurations.
@@ 4.1.1 Scenario A: Control Performance of Jet Pair
In this study, we use DRL to develop flow control strategies for the flow around a square cylinder, as shown in figure 14. The figure provides a detailed analysis of the influence of synthetic jet positioning on flow control performance within the square geometry. The reward function serves as the primary metric for evaluating the effectiveness of the control strategies developed by the DRL agent. Initially, the reward values are low, reflecting the early stages of learning and the exploration of suboptimal strategies. As training progresses, a significant increase in the reward function is observed across all jet configurations, particularly between the 500th and 2000th episodes. This trend indicates that the DRL agent is successfully optimizing the control strategies, leading to improved flow control. The consistent rise in the reward function suggests that the DRL algorithm is steadily refining its policy, converging toward an optimal solution. The agent’s ability to adapt and improve through continuous learning highlights the potential of DRL in addressing complex flow control challenges.
@@
We examined the reward function behavior during DRL training with synthetic jets positioned near the front and rear corners of a square cylinder, specifically for the configurations $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$, as illustrated in figure 14(a). Our findings reveal that the reward function experienced the most rapid growth when the jets were located at $J_{3}^{1},J_{3}^{2}$, whereas the slowest growth occurred for the $J_{2}^{1},J_{2}^{2}$ configuration. This variation in growth rates underscores the superior adaptability of the $J_{3}^{1},J_{3}^{2}$ configuration, as it allows the DRL agent to more effectively achieve the control objectives. The accelerated increase in the reward function associated with $J_{3}^{1},J_{3}^{2}$ indicates that this setup is intrinsically more aligned with the flow control requirements, facilitating a more efficient optimization process.
@@
In contrast, the slower growth observed for $J_{2}^{1},J_{2}^{2}$ suggests that this configuration may encounter challenges in interacting with the flow dynamics, potentially due to a more complex flow behavior that hinders the learning efficiency of the agent. These findings highlight the critical role of jet positioning in optimizing flow control strategies. The distinct differences in reward growth rates not only reflect the varying degrees of learning efficiency but also emphasize the importance of strategic jet placement in enhancing the performance of DRL-based flow control. Configurations like $J_{3}^{1},J_{3}^{2}$, which demonstrate rapid reward growth, are evidently more conducive to achieving efficient and precise flow manipulation, whereas slower growth configurations, such as $J_{2}^{1},J_{2}^{2}$, may require more refined strategies or adjustments to overcome inherent complexities. These insights are vital for refining jet configurations and advancing the effectiveness of DRL algorithms in complex fluid dynamics applications.
@@
We present the relationship between the drag coefficient and non-dimensional time for four different jet placement configurations, compared to the baseline case without active control, as shown in figure 14(b). The baseline curve, representing the drag coefficient without control, exhibits periodic oscillations. Once the synthetic jets are activated, the drag coefficient of the square cylinder decreases significantly across all configurations. The most notable reduction occurs when the jets are positioned near the rear corner points at $J_{3}^{1},J_{3}^{2}$, indicating this placement is the most effective for drag reduction. In contrast, jets placed near the front corners at $J_{1}^{1},J_{1}^{2}$ show a clear reduction in drag, but with transient spikes and oscillations, suggesting a dynamic interaction between the jets and the flow. Configurations at $J_{2}^{1},J_{2}^{2}$ and $J_{4}^{1},J_{4}^{2}$ also result in substantial drag reduction, though less pronounced than at $J_{3}^{1},J_{3}^{2}$. This analysis highlights the critical role of jet placement in optimizing flow control and reducing aerodynamic drag around square cylinders.
@@
We depict the relationship between the lift coefficient and non-dimensional time for various synthetic jets configurations, compared to the baseline scenario, as shown in figure 14(c). The baseline lift coefficient exhibits periodic fluctuations throughout the time series, characteristic of the natural vortex shedding around the square cylinder. When the synthetic jets are positioned at $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$, a significant attenuation in the lift coefficient is observed. Specifically, the configuration at $J_{3}^{1},J_{3}^{2}$ induces the most pronounced initial reduction in the lift coefficient, which rapidly stabilizes near zero, indicating effective suppression of the unsteady aerodynamic forces. Furthermore, the configurations at $J_{2}^{1},J_{2}^{2}$ and $J_{4}^{1},J_{4}^{2}$ lead to a marked reduction in both the amplitude and frequency of the lift coefficient oscillations, thereby significantly damping the lift fluctuations. This suggests a successful mitigation of vortex-induced forces through controlled jet actuation. However, when the jets are located at $J_{1}^{1},J_{1}^{2}$, the lift coefficient fails to exhibit a similar reduction. Instead, transient spikes and persistent oscillations are present, indicating that this jet configuration is less effective in altering the flow structures responsible for lift generation. These results crystallize the crucial role of optimal jet placement in flow control, where precise modulation of aerodynamic forces, such as lift, is essential for achieving effective performance. Jet configurations that interact favorably with the complex flow dynamics around the square cylinder are the most effective in stabilizing the flow and reducing oscillatory forces.
4.1.2 Scenario A: Energy Consumption of Jet Pair
@@
Figure 15: Energy consumption over non-dimensional time for jet pairs positioned at various locations around the square cylinder. The jet pair locations include: $J_{1}^{1},J_{1}^{2}$, $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$.
@@
The external energy consumed for flow control varies significantly with the positioning of the synthetic jets at $J_{1}^{1},J_{1}^{2}$, $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$. The time evolution of energy consumption for these different configurations is illustrated in figure 15. When the jets are located at $J_{1}^{1},J_{1}^{2}$, the energy required for flow control is substantially higher, accompanied by pronounced fluctuations, indicating instability in mass flow rates. This instability suggests lower control efficiency due to complex interactions with the flow field, making it difficult to achieve consistent and effective flow manipulation. In contrast, when the jets are positioned at $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$, the control strategies exhibit more stable behavior, with minimal oscillations and near-zero values, reflecting efficient and balanced flow control with lower energy input. Notably, the $J_{3}^{1},J_{3}^{2}$ configuration requires the least energy at the onset of control and remains the most stable over time. The comparison between the less stable $J_{1}^{1},J_{1}^{2}$ configuration and the more stable configurations highlights the differences in energy efficiency and control stability. The larger and more unstable fluctuations observed with $J_{1}^{1},J_{1}^{2}$ indicate lower energy efficiency and greater challenges in maintaining flow control. In contrast, $J_{3}^{1},J_{3}^{2}$ demonstrates more predictable and controlled energy usage, effectively sustaining flow control with lower energy costs.
@@
In figure 16, the activation states of the synthetic jets at positions $J_{1}$, $J_{2}$, $J_{3}$, and $J_{4}$ are depicted, along with the velocity contours of the flow field after flow control is completed. The velocity contours show that when the jets are positioned near the rear sides of the square cylinder at $J_{3}$, vortex shedding in the wake is fully suppressed. At $J_{4}$, located on the rear face of the cylinder, the recirculation region expands significantly compared to the initial stages of control. When the jets are placed closer to the upstream at $J_{2}$, periodic actuation of the synthetic jets enhances the flow stability around the cylinder. Finally, at $J_{1}$, positioned at the front of the cylinder, the interaction between the synthetic jets and the free stream improves flow stability, although alternating vortex shedding still persists.
Overall, the control performance of the jets varies significantly with placement. When the synthetic jets are positioned near the front corner points, vortex shedding is not fully suppressed due to the large separated flow reattaching along the sides of the square cylinder before separating again near the rear corners. In contrast, placing the jets near the rear corner points at position $J_{3}$ provides more effective flow control, leading to the complete suppression of vortex shedding. Chen et al. (2023) used DRL-based flow control to reduce vortex-induced vibrations around a square cylinder ($Re=100$) and compared the control effects of jet placement at the front, middle, and rear positions along the cylinder’s side. Their results showed that full suppression was achieved only near the rear corners, consistent with our findings.

### v5

@@ 3.2.2 Scenario B: Energy Consumption with Simultaneous Control (Two Pairs)
Figure 11: Comparison of action values over non-dimensional time for different synthetic jets configurations around the cylinder. (a) Two pairs of jets: $J_{2}^{1},J_{2}^{2}$ and $J_{3}^{1},J_{3}^{2}$. (b) Two pairs of jets: $J_{3}^{1},J_{3}^{2}$ and $J_{4}^{1},J_{4}^{2}$.
A flow control strategy involving the mass flow rates of two pairs of synthetic jets is implemented, which is referred to as “action” in figure 11. The action values represent the external energy consumed to execute flow control, and the evolution of energy consumption for various jet configurations is plotted over dimensionless time. Figure 11(a) illustrates the energy consumption of the control strategy utilizing two pairs of synthetic jets: $J_{3}^{1},J_{3}^{2}$ and $J_{2}^{1},J_{2}^{2}$. The initial control actions display significant peaks, reflecting the high energy required at the onset of the flow control process. Over time, both jet configurations show a decrease in energy consumption, with the control efforts stabilizing. However, the jet pair $J_{3}^{1},J_{3}^{2}$ exhibits a slightly higher energy demand compared to the pair $J_{2}^{1},J_{2}^{2}$. This indicates that the jets positioned at $J_{2}^{1},J_{2}^{2}$ may be more energy-efficient in maintaining effective flow control.
Figure 11(b) illustrates the control strategy using two pairs of synthetic jets, namely $J_{3}^{1},J_{3}^{2}$ and $J_{4}^{1},J_{4}^{2}$. The initial peak in the action values signifies the maximum energy required at the onset of the flow control process. Over time, the energy consumption stabilizes, indicating that the control strategy becomes more consistent. In the long term, the jet pair $J_{4}^{1},J_{4}^{2}$ consumes slightly less energy compared to the jet pair $J_{3}^{1},J_{3}^{2}$, suggesting that positioning the synthetic jets at $J_{4}$ may be more energy-efficient than jets at $J_{3}$. In summary, the initial peak in the control strategy underscores the substantial energy demand at the inception of the process, while the subsequent stabilization reflects the sustained maintenance of flow control. Moreover, the disparities in energy consumption across different configurations provide valuable insights for optimizing flow control efficiency through judicious selection of jet placement. Specifically, jets positioned at $J_{2}$ locations exhibit superior energy efficiency compared to those at $J_{3}$ locations, and similarly, $J_{4}$ locations demonstrate enhanced energy savings relative to $J_{3}$ locations.
@@ 3.2.3 Scenario B: Energy Consumption with Simultaneous Control (Three Pairs)
Figure 12 presents a detailed analysis of energy consumption for the control strategy involving the simultaneous activation of three pairs of synthetic jets. The jets—$J_{2}^{1}$ and $J_{2}^{2}$, $J_{3}^{1}$ and $J_{3}^{2}$, and $J_{4}^{1}$ and $J_{4}^{2}$—are symmetrically positioned around the cylinder at angular locations of 105°, 90°, and 75° from the forward stagnation point. The external energy required for the synthetic jet pair $J_{3}$ (comprised of $J^{1}_{3}$ and $J^{2}_{3}$) is denoted by $E_{J_{3}}$, which is expressed as the integral of the control actions over time. Interestingly, when compared to figure 11, where only two pairs of synthetic jets are employed, figure 12 demonstrates that the activation of three pairs of synthetic jets leads to a reduction in the peak external energy demand for each individual jet. This suggests that the simultaneous use of additional jets can distribute the energy load more effectively, thereby enhancing the overall efficiency of the flow control strategy.
@@
Figure 12: Energy consumption analysis for different jet pairs positioned around a cylinder. The left image shows the jet pairs $J_{2}$, $J_{3}$, and $J_{4}$, with azimuth angles ranging from 75° to 105°. The right image presents the corresponding actuation values of the synthetic jets over non-dimensional time for each jet configuration.
All three jet configurations exhibit pronounced peaks in the DRL-optimized control actions during the initial transient phase, followed by a stabilization of energy consumption over time. Among them, the configuration with jets located at $J_{3}$ shows the highest peak in control effort. The time-integrated control actions quantify the total energy input required for effective flow regulation, revealing that the energy consumption at $E_{J_{3}}$ is significantly higher than that of the other two configurations, with $E_{J_{4}}\approx E_{J_{2}}$. This observation suggests that the 90° jet placement demands greater energy input to achieve comparable control performance, indicating reduced energy efficiency relative to the $J_{4}$ and $J_{2}$ configurations. Moreover, between the latter two, jets at $J_{2}$ require less total energy than those at $J_{4}$, further highlighting $J_{2}$ as the most energy-efficient option. Figure 12 summarizes the energy expenditure associated with different jet placements, emphasizing the importance of actuator positioning in achieving energy-efficient yet effective flow control.
Based on the analysis of DRL-based control strategies involving multiple pairs of synthetic jets, several key advantages over traditional flow control approaches are identified. One notable feature is that, regardless of whether two or three jet pairs are employed, no prior specification of actuation parameters such as frequency or amplitude is required. The control strategies are adaptively learned by the DRL agent through direct interaction with the flow field, eliminating the need for manual tuning. This highlights a fundamental distinction from conventional methods. In comparison to traditional approaches, the DRL-based framework offers three major advantages. First, it removes the requirement for predefined control parameters, allowing the agent to autonomously infer effective strategies from flow feedback, thereby avoiding the exhaustive parameter studies typically required in periodic control schemes, such as those implemented by Li et al.. Second, it achieves superior control performance; while conventional techniques—e.g., those using control rods as studied by Lu et al.—may reduce drag or lift to some extent, they generally fail to fully suppress vortex shedding. In contrast, the DRL-based controller can stabilize the wake entirely. Third, unlike genetic algorithms such as that employed by Zigunov et al., which operate within predefined parameter spaces and risk missing global optima, DRL enables adaptive global exploration without prior knowledge of control laws. This results in significantly improved optimization efficiency and broader applicability to complex flow scenarios.
@@ 4.1 A pair of synthetic jets controllers
4.1.1 Scenario A: Control Performance of Jet Pairs
@@
DRL is used to develop flow control strategies for the flow around a square cylinder, as shown in figure 13. The figure provides a detailed analysis of the influence of synthetic jet positioning on flow control performance within the square geometry. The reward function serves as the primary metric for evaluating the effectiveness of the control strategies developed by the DRL agent. Initially, the reward values are low, reflecting the early stages of learning and the exploration of suboptimal strategies. As training progresses, a significant increase in the reward function is observed across all jet configurations, particularly between the 500 and 2000 episodes. This trend indicates that the DRL agent is successfully optimizing the control strategies, leading to improved flow control. The consistent rise in the reward function suggests that the DRL algorithm is steadily refining its policy, converging toward an optimal solution.
Figure 13: Performance evaluation of different jet configurations for flow control around a square cylinder. (a) reward curves over training episodes. (b) variation of $C_{D}$ over time, compared with the baseline scenario. (c) variation of $C_{L}$ over time, compared with the baseline scenario.
@@
The agent’s ability to adapt and improve through continuous learning highlights the potential of DRL in addressing complex flow control challenges. The reward function behavior during DRL training is examined with synthetic jets positioned near the front and rear corners of a square cylinder, specifically for the configurations $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$, as illustrated in figure 13(a). The results indicate that the reward function exhibited the most rapid growth when the jets are positioned at $J_{3}^{1},J_{3}^{2}$, while the slowest growth was observed for the $J_{2}^{1},J_{2}^{2}$ configuration. This variation in growth rates underscores the superior adaptability of the $J_{3}^{1},J_{3}^{2}$ configuration, as it allows the DRL agent to more effectively achieve the control objectives. The accelerated increase in the reward function associated with $J_{3}^{1},J_{3}^{2}$ indicates that this setup is intrinsically more aligned with the flow control requirements, facilitating a more efficient optimization process.
Table 1: Control effectiveness and energy consumption analysis using a single pair of synthetic jets on a square cylinder.
Jet $C_{\text{D,Base}}$ $C_{\text{D,Mean}}$ Reduction (%) $C_{\text{D,Std}}$ $C_{\text{L,Mean}}$ $C_{\text{L,Std}}$ $A_{\text{Std}}$ $A_{\text{Mean}}$ Action ratio (%)
$J_{1}$ 1.549 1.361 12.1 0.028 0.008 0.164 0.880 19.594 44.0
$J_{2}$ 1.549 1.344 13.2 0.002 0.072 0.030 -0.506 0.177 25.3
$J_{3}$ 1.549 1.326 14.4 0.002 0.008 0.013 -0.033 0.037 1.7
$J_{4}$ 1.549 1.353 12.6 0.001 -0.012 0.043 0.184 0.654 9.2
@@
In contrast, the slower growth observed for $J_{2}^{1},J_{2}^{2}$ suggests that this configuration may encounter challenges in interacting with the flow dynamics, potentially due to a more complex flow behavior that hinders the learning efficiency of the agent. These findings highlight the critical role of jet positioning in optimizing flow control strategies. The distinct differences in reward growth rates not only reflect the varying degrees of learning efficiency but also emphasize the importance of strategic jet placement in enhancing the performance of DRL-based flow control. Configurations like $J_{3}^{1},J_{3}^{2}$, which demonstrate rapid reward growth, are evidently more conducive to achieving efficient and precise flow manipulation, whereas slower growth configurations, such as $J_{2}^{1},J_{2}^{2}$, may require more refined strategies or adjustments to overcome inherent complexities.
@@
The relationship between the drag coefficient and non-dimensional time is presented for four different jet placement configurations, compared to the baseline case without active control, as shown in figure 13(b). The baseline curve, representing the drag coefficient without control, exhibits periodic oscillations. Once the synthetic jets are activated, the drag coefficient of the square cylinder decreases significantly across all configurations. The most notable reduction occurs when the jets are positioned near the rear corner points at $J_{3}^{1},J_{3}^{2}$, indicating this placement is the most effective for drag reduction. In contrast, jets placed near the front corners at $J_{1}^{1},J_{1}^{2}$ show a clear reduction in drag, but with transient spikes and oscillations, suggesting a dynamic interaction between the jets and the flow. Configurations at $J_{2}^{1},J_{2}^{2}$ and $J_{4}^{1},J_{4}^{2}$ also result in substantial drag reduction, though less pronounced than at $J_{3}^{1},J_{3}^{2}$.
@@
The relationship between the lift coefficient and non-dimensional time is depicted for various synthetic jet configurations, compared to the baseline scenario, as shown in figure 13(c). The baseline lift coefficient exhibits periodic fluctuations throughout the time series, characteristic of the natural vortex shedding around the square cylinder. When the synthetic jets are positioned at $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$, a significant attenuation in the lift coefficient is observed. Specifically, the configuration at $J_{3}^{1},J_{3}^{2}$ induces the most pronounced initial reduction in the lift coefficient, which rapidly stabilizes near zero, indicating effective suppression of the unsteady aerodynamic forces. Furthermore, the configurations at $J_{2}^{1},J_{2}^{2}$ and $J_{4}^{1},J_{4}^{2}$ lead to a marked reduction in both the amplitude and frequency of the lift coefficient oscillations, thereby significantly damping the lift fluctuations. This suggests a successful mitigation of vortex-induced forces through controlled jet actuation. However, when the jets are located at $J_{1}^{1},J_{1}^{2}$, the lift coefficient fails to exhibit a similar reduction. Instead, transient spikes and persistent oscillations are present, indicating that this jet configuration is less effective in altering the flow structures responsible for lift generation. These results crystallize the crucial role of optimal jet placement in flow control, where precise modulation of aerodynamic forces, such as lift, is essential for achieving effective performance.
4.1.2 Scenario A: Energy Consumption of Jet Pairs
@@
Figure 14: (a) Energy consumption over non-dimensional time for jet pairs positioned at various locations around the square cylinder. (b) Streamlines of the controlled flow and the baseline flow when the synthetic jet is located at different positions.
Figure 14 presents a comparative visualization of the synthetic jet actuation history and corresponding flow structures for selected jet configurations. Figure 14 (a) illustrates the temporal evolution of the control action, represented as the normalized mass flow rate of synthetic jets, for four distinct jet placements denoted as $J_{1}^{1}J_{1}^{2}$, $J_{2}^{1}J_{2}^{2}$, $J_{3}^{1}J_{3}^{2}$, and $J_{4}^{1}J_{4}^{2}$. The control input, expressed as the non-dimensionalized mass flow rate of the synthetic jet, is actively modulated by the reinforcement learning agent. All control signals exhibit oscillatory behavior with zero mean, a typical feature of synthetic jets. However, the amplitude, frequency, and structure of the control inputs vary significantly depending on the jet placement.
@@
The external energy consumed for flow control varies significantly with the locations of the synthetic jets at $J_{1}^{1},J_{1}^{2}$, $J_{2}^{1},J_{2}^{2}$, $J_{3}^{1},J_{3}^{2}$, and $J_{4}^{1},J_{4}^{2}$. When the jets are positioned at $J_{1}^{1}$ and $J_{1}^{2}$, the energy required for flow control increases markedly, accompanied by pronounced fluctuations, indicating unstable mass flow rates. This instability suggests that the complex interactions with the flow field result in lower control efficiency, making consistent and effective flow manipulation challenging. In contrast, when the jets are located at the other positions, the control strategy exhibits more stable behavior with minimal oscillations and values close to zero, reflecting efficient and balanced flow control with lower energy input. Specifically, the fluctuation amplitudes of the control actions at $J_{2}^{1},J_{2}^{2}$ and $J_{4}^{1},J_{4}^{2}$ are notably smaller than those at $J_{1}^{1},J_{1}^{2}$. Notably, the configuration with jets at $J_{3}^{1}$ and $J_{3}^{2}$ requires the least energy at the onset of control and demonstrates the most stable control actions over time. The comparison between the less stable $J_{1}^{1},J_{1}^{2}$ configuration and the more stable $J_{3}^{1},J_{3}^{2}$ configuration highlights the differences in energy efficiency and control stability.
The differences in the observed control signals correspond to distinct flow modifications illustrated in figure 14(a), where the streamlines reveal varying degrees of vortex suppression. The configuration with the highest energy consumption, $J_{1}^{1}J_{1}^{2}$, results in the least effective flow control, as shown in figure 14(c). Moderate energy-consuming placements, $J_{2}^{1}J_{2}^{2}$ and $J_{4}^{1}J_{4}^{2}$, yield improved wake stability compared to the baseline case, as depicted in figures 14(d) and 14(e), respectively. The most energy-efficient configuration, $J_{3}^{1}J_{3}^{2}$, leads to the most stabilized wake structure, as shown in figure 14(f), where the wake length is significantly shortened and the vortex intensity is markedly reduced, demonstrating superior control performance. In contrast, other jet placements induce less pronounced modifications in the wake. These findings highlight the critical importance of jet placement in determining the effectiveness of synthetic jet-based flow control strategies.
@@
Figure 15 presents the activation states of the synthetic jets at positions $J_{1}$, $J_{2}$, $J_{3}$, and $J_{4}$, alongside the corresponding velocity contours of the controlled flow field. The left column illustrates the instantaneous velocity fields with active jet actuation, while the right column depicts the resulting wake structures. The velocity contours reveal that when the synthetic jets are positioned near the rear corners of the square cylinder at $J_{3}$, vortex shedding is entirely suppressed, yielding a highly stabilized wake. The synthetic jets at $J_{4}$, located on the rear face of the cylinder, the recirculation region expands significantly compared to the initial stages of control. The synthetic jets at $J_{2}$, where the jets are located near the front corners, periodic actuation interacts with the shear layers, mitigating flow separation and enhancing wake. Finally, when the jets are positioned at $J_{1}$ on the front face, their interaction with the incoming free stream improves near-wake stability; however, alternating vortex shedding persists, suggesting that upstream jet placement is less effective in fully eliminating wake instabilities. These results underscore the critical role of jet positioning in flow control efficacy, demonstrating that optimal placement at $J_{3}$ achieves complete vortex suppression, whereas other configurations yield varying degrees of wake stabilization and turbulence modulation.
The control performance and energy consumption of flow control based on DRL are summarized in table 1. The control performance is evaluated by the percentage reduction in drag coefficient $C_{\text{D,Mean}}$ relative to the baseline drag coefficient $C_{\text{D,Base}}$, the standard deviation of the drag coefficient $C_{\text{D,Std}}$ for the controlled flow, and the mean $C_{\text{L,Mean}}$ and standard deviation $C_{\text{L,Std}}$ of the lift coefficient. The energy consumption for flow control is quantified by the mean $A_{\text{Mean}}$ and standard deviation $A_{\text{Std}}$ of the jet actuation magnitude, as well as the action ratio, which is the percentage of the average jet flow rate relative to the upstream inlet flow rate. The quantified control performance and energy consumption metrics in the table indicate that positioning the synthetic jet at $J_{3}$ achieves the highest drag reduction efficiency of 14.4%, using only 1.7% of the inlet flow rate, while also minimizing the standard deviations of $C_{D}$, $C_{L}$, and jet actuation. In contrast, placing the jet at $J_{1}$ requires 44% of the inlet flow rate to achieve a drag reduction of only 12.1%. This configuration also exhibits the highest standard deviations in $C_{D}$, $C_{L}$, and jet actuation, indicating the least stability.
Overall, the control performance of the jets varies significantly with placement. When the synthetic jets are positioned near the front corner points, vortex shedding is not fully suppressed due to the large separated flow reattaching along the sides of the square cylinder before separating again near the rear corners. In contrast, placing the jets near the rear corner points at position $J_{3}$ provides more effective flow control, leading to the complete suppression of vortex shedding. Chen et al. used DRL-based flow control to reduce vortex-induced vibrations around a square cylinder and compared the control effects of jet placement at the front, middle, and rear positions along the cylinder’s side. Their results showed that full suppression was achieved only near the rear corners, consistent with our findings Chen et al. (2023).